OpenStudy (anonymous):

Grant can row 6 miles in 36 minutes. Arianna can row 3,520 yards in 20 minutes. In miles per hour, how much faster does Grant row than Arianna?

OpenStudy (anonymous):

Isn't it like 1500 yds is equal to a mile?

OpenStudy (anonymous):

1 mile = 5280 feet. yds = 5280 / 3

OpenStudy (anonymous):

1mile = 1760 yds. 6miles = 10560 yds. 10560/ 36 = yds/min = 3520/20 = yds/ min

OpenStudy (anonymous):

i need help setting it up ion get it

OpenStudy (anonymous):

Convert it to like terms. What don't you understand, I'd be happy to explain.

OpenStudy (anonymous):

i dont know how to find the answer

OpenStudy (anonymous):

I just went through how to do it in my above descriptions.

OpenStudy (anonymous):

dont know how to do that

OpenStudy (anonymous):

anybody know the answer will give medal!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

I literally gave you the answer. Grant: 10560/ 36 = yds/min = Arianna: 3520/20 = yds/ min

OpenStudy (anonymous):

Just get a calculator and do those divisions. You're here to learn. Not to be given homework answers.

OpenStudy (anonymous):

for the first one i got 29.3 and for the second one i got 176